OpenStudy (anonymous):

Hi... I need help to check my answers. Elimination or Addition. 2x+3y+z=7 x-2y+z=-4 -x+3y-z=6

OpenStudy (anonymous):

@waterineyes my answers are x=1 z=-1 y=2

OpenStudy (anonymous):

1,2-1

OpenStudy (anonymous):

Yes you are absolutely right dear.. \[\large \color{green}{\text{Keep It Up..}} \color{red} {\checkmark}\]

OpenStudy (anonymous):

Thanks a lot dear. Gonna solve one problem and please help me check...

OpenStudy (anonymous):

Sure.. Try it on your own first..

OpenStudy (anonymous):

yes I will dear. Thanks a lot!

OpenStudy (anonymous):

if you ever in doubt of your answer just plug it in to one of the equation :)

OpenStudy (anonymous):

Thanks @timo86m

OpenStudy (anonymous):

2x+3y+z=7 1,2,-1=x,y,z 2*1+3*2+(-1) = 7 2+6-1=7 8-1=7 7=7

OpenStudy (anonymous):

Thanks for helping me how to check @timo86m

OpenStudy (anonymous):

De nada. And thanks for the medal :D

OpenStudy (anonymous):

I give you a medal for that @timo86m

OpenStudy (anonymous):

@nia_maepenix yesterday I was pointing to my post.. If you read that post then you could know that to verify our answers are right we simply put them in the equations if they come equal then our answers are right..

OpenStudy (anonymous):

Thanks to you @waterineyes I did good in my math yesterday!